[Brahms-l] Shift report 20060619 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Jun 19 2006 - 00:01:54 EDT
Shifter leader was SB
Shifters were 


1. Beam from day shift continued into the evening shift.
   Beam has been stable for the whole shift, continues into
   the night shift

2. Runs @2.3 deg @1723 A 
   16231 - 16240, 16241 still going on

3. No new detector problems

   For some unknown reason, information on 
      Specie 
      momentum 
      fill number 

   are not recoreded in the E-log for run 16240. 
   The data seems fine looking at supermon. 
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 19 00:02:22 2006

This archive was generated by hypermail 2.1.8 : Mon Jun 19 2006 - 00:02:37 EDT